How many different 9 player baseball teams can be formed from 15 players if 3 only pitch while the remaining 12 can play any of the 8 remaining positions?
I really do not have a clue how to solve. I tried to use different types of combination things that i had learned in class, but it has too many variables for it.

Found 2 solutions by stanbon, mathslover:
Answer by stanbon(75887)

How many different 9 player baseball teams can be formed from 15 players if 3 only pitch while the remaining 12 can play any of the 8 remaining positions?
I really do not have a clue how to solve.
---------
Pick a pitcher in 3 ways
Pick the remaining 8 players: 12C8 = 12C4 = 495 ways
-----------
# of teams = 3*495 = 1485
======================
Cheers,
Stan H.

Answer by mathslover(157)

you can select a pitcher among the three available in
3C1 way. the remaining 8 have to be selected from 12 players
number of ways of selecting the remaining 8 players is 12C8
total combination of teams possible = 3C1 * 12C8
= 3 * (12*11*10*9)/(4*3*2*1)
=1350